a boat can maintain a constant speed of 50mil/hr relative to the water. the boat makes a trip up stream in 57min; the return trip takes 43min. what is the speed of the current?
Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

57/60(50-c)=43/60(50+c)
57(50-c)=43(50+c)
2850-57c=2150+43c
700=100c
c=7
The current is 7 mph..

Answer by Alan3354(69443)

a boat can maintain a constant speed of 50mil/hr relative to the water. the boat makes a trip up stream in 57min; the return trip takes 43min. what is the speed of the current?
--------------------
The general solution is:
c = v*(diff in t)/(sum of t)
c = 50*(57 - 43)/(57 + 43) = 50*14/100
c = 7 mph